OpenStudy (emilyjones284):

Find absolute value V and absolute value U+V, given that u=-16j and v=-5i-2j

OpenStudy (amigatour):

The value of: U=-16j V=-5i-2j What's your question? You should ask questions. It sounds that you're testing people. LOL.

OpenStudy (emilyjones284):

The part where I said Find the... is my question haha. I dont know how to find it

ganeshie8 (ganeshie8):

absolute value of \(ai+bj\) is \(\sqrt{a^2 + b^2}\)

ganeshie8 (ganeshie8):

\(v = -5i - 2j\) absolute value = \(\sqrt{(-5)^2 + (-2)^2}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (emilyjones284):

so would I leave my answer as - square root of 29?

OpenStudy (emilyjones284):

and how would I find absolute value of U+V? Thats the part I'm having trouble with

ganeshie8 (ganeshie8):

yes absolute value of v is sqrt(29)

ganeshie8 (ganeshie8):

first find U+V

ganeshie8 (ganeshie8):

\(u=-16j\) \( v=-5i-2j \)

ganeshie8 (ganeshie8):

\(u + v = -16j - 5i - 2j\)

ganeshie8 (ganeshie8):

simplify

ganeshie8 (ganeshie8):

add like terms

ganeshie8 (ganeshie8):

\(u + v = - 5i - 18j\)

ganeshie8 (ganeshie8):

right ?

OpenStudy (emilyjones284):

Yup thats what I got

ganeshie8 (ganeshie8):

take the absolute value of that

OpenStudy (emilyjones284):

So I don't have to do anything else? I remember it being more complex in my math class

OpenStudy (emilyjones284):

Ohhh

ganeshie8 (ganeshie8):

absolute value of \(-5i - 18j\) is ?

OpenStudy (emilyjones284):

5i+18j?

ganeshie8 (ganeshie8):

no, this absolute value is different.

ganeshie8 (ganeshie8):

in vectors, absolute value means "distance"

ganeshie8 (ganeshie8):

absolute value of \(-5i - 18j\) is : \(\sqrt{(-5)^2 + (-18)^2}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (emilyjones284):

Ohhh sorry I got it now

ganeshie8 (ganeshie8):

good :)

OpenStudy (emilyjones284):

So square root of 349?

ganeshie8 (ganeshie8):

sqrt(349) is \(\large \color{red}{\checkmark}\)

OpenStudy (emilyjones284):

Cool! So that's all I need to do to get the answer?

ganeshie8 (ganeshie8):

yes !

OpenStudy (emilyjones284):

Yay! Thank you!